The x intercept of the graph of 2x+3y=5 is

Please show all work
Thank You

Q23

Answer by chitra(359)

The x intercept of the graph of 2x + 3y = 5

To find the x intercept of any equation plug in y = 0 in any equation.

So when y = 0 in the given equation, gives us

2x + 3(0) = 5

2x + 0 = 5

2x = 5

x = 5/2

thus, the solution